A convex surface has a uniform radius of curvature ‘5R’. A wheel of radius ‘R’ is rolling without slipping on it with its centre of mass moving with constant speed ‘v’. The acceleration of the point ‘P’ in contact with convex surface is

Grade:11

2 Answers

Arun
5 years ago
 
Dear student
 
 
centrifugal force F = mv2/r
acceleration = F/m = v2/r = v2/5R
 
 
 
Regards
Arun (askIITians forum expert)
ARYAN CHORDIA
3 years ago
acceleration of point=acceleration with respect to centre of mass +acceleration of centre of mass
acceleration of COM=v^2/(5R+R)=v^2/6R
acceleration of P wrt com=0 as the point has zero velocity due to pure rolling
Therefore total accln=v^2/6R